# Correlations Are Ruining Your Gradient Descent

### v3

@@ 5 Discussion
5.1 Decorrelation enables approximations of gradient descent
The success of decorrelation in enabling learning with the node perturbation and feedback alignment algorithms indicates that, via removal of cross-correlation structure between nodes in layers of a neural network, minimisation of the loss function is made much ‘easier’. These two algorithms are by no means the only algorithms which might be enabled by a decorrelation mechanism, and if one is interested in training of neural networks upon exotic hardware, one might consider combining such a distributed decorrelation mechanism with a number of alternative learning algorithms such as direct feedback alignment [Lillicrap et al., 2016, Nøkland, 2016], surrogate gradient learning Neftci et al. [2019] and more.
5.2 Decorrelation in biological systems
Yet another parallel stream of this work which has engaged with concepts of decorrelation and whitening is that which has taken place in the context of biological neural networks, both on a computational and experimental level.
@@ 5.2 Decorrelation in biological systems
In the neurosciences, decorrelation has been presented as an explanation of the filtering which happens at multiple levels in nervous systems. The center-surround processing which takes places at the earliest stage of visual processing (at retinal ganglion cells), for example, has been proposed as an aid for decorrelation of visual input to the brain [Pitkow and Meister, 2012]. At a more general scale, inhibitory plasticity appears to be learned in a manner which also leads to spatial decorrelation [He and Cline, 2019], something which has also been modelled [Vogels et al., 2011].
@@
Beyond experimental work, simulation work has also shown that via decorrelation of unit activities, competitive learning can be established to learn subspaces in a distributed and ‘local’ fashion [Bell and Sejnowski, 1997, Földiák, 1990, Oja, 1989], something which can also be extended for the reproduction of early visual cortical receptive fields [Zylberberg et al., 2011].
Though much of this work is framed by work in computational neuroscience, the strict decorrelation which we employ to overcome the issues inherent to gradient descent is far from the kind of decorrelation observed in real neural systems. Neural data suggests that nearby neurons can have correlated activities - a feature which may be required for redundancy and robustness to cell death. Furthermore, the decorrelation mechanisms which have been observed are often learned in a slow manner and on many different scales depending upon the brain area under consideration. All in all, though we point toward a promising method by which local and distributed learning can be enabled, there remain significant questions about how much of this analysis might be directly mappable to neural systems. Finally, a more detailed model of neuronal interaction, for example in the style proposed by Lipshutz and Simoncelli [2024], would be more appropriate for providing a detailed proposal on the role of different neuron types in decorrelation.
5.3 Decorrelation and overfitting
Wadia et al. [2021] pointed out that the restriction to whitened data or second-order optimization methods can restrict the space of generalization unless properly regularized. This issue is one to keep in mind when developing methods which speed up training of models significantly and generalization performance should be a point of concern. However, as also described by Wadia et al. [2021], it may be that a regularized form of whitening/decorrelation would in fact be optimal for training for maximum generalization performance.
We find that our decorrelation approach has a combination of benefits: increased convergence speed along with increased generalization performance. One core difference of our approach is a focus upon decorrelation over whitening. Whitening focuses upon not only decorrelation but a rescaling of all unit activities for a response variance of one. We speculate that this causes a rescaling of not only ‘signal’ inherent in network activities, but also ‘noise’, and that noise is inherently of a smaller scale than signal in trained multi-layer networks. Thus we propose that decorrelation might be a better tradeoff, where signal correlations are removed without rescaling them and thus the scale of noise (or training-data specific signal variations) is not increased and used for overfitting.
@@ 6 Conclusion
Here we present an integration of a range of research directions including natural gradient descent, decorrelation and whitening, as well as approximate or alternative methods for gradient descent. Notably we illustrate and describe how one component of natural gradients is often overlooked and can be framed as correlations in feature data (at all layers of a neural network) bringing parameters into a non-orthonormal relationship. We show that data correlation in neural networks can be removed, in a similar fashion to neural competition, through an additional transformation at each layer of a network. Removing these correlations allows much faster training of neural networks with backpropagation but also magnitudes faster training of neural networks via alternative approximate algorithms. These results and insights together suggest that failures of ‘biologically-plausible’ learning approaches and learning rules for distributed computing could be overcome through decorrelation and the return of parameters to an orthogonal basis.
@@ Appendix C A note on decorrelation vs whitening
In this work we limit our emphasis on whitening and focus more upon decorrelation. The reason for this is two-fold. First, the off-diagonal elements of the decorrelation matrices cause the greatest impact in skewing gradient descent. The diagonal elements simply scale up/down the gradient vector and this can be trivially dealt with by modern optimizers (e.g. with the Adam optimizer [Kingma and Ba, 2014]). Second, for data in which features are often zero or extremely sparse, normalizing for unit variance can result in extremely large valued activations (distributions with extremely long tails) or unstable updates in the stochastic updating regime.
@@ Appendix E Training pseudocode
Algorithm 1 describes pseudocode for the full computation and updating of a network which is carried out in this work. This is shown assuming the presentation of a single sample with its corresponding network update, though is in-practice used with a mini-batch of size 256 by default. The algorithm is shown for the case of backpropagation, though it is compatible with alternative credit assignment methods such as feedback alignment or node perturbation. Furthermore, the pseudocode shows the forward parameters as being updated by SGD, though in practice these are updated with the Adam [Kingma and Ba, 2014] optimizer.

### v4

@@ 5 Discussion
Herein we were able to link the concepts of natural gradients and decorrelation to show how, why, and to what degree decorrelating node activities at every layer of a neural network can massively boost performance of approximate gradient descent methods. The two approximate gradient descent algorithms treated herein are by no means the only algorithms which might be enabled by a decorrelation mechanism, and if one is interested in training of neural networks upon exotic hardware, one might consider combining such a distributed decorrelation mechanism with a number of alternative learning algorithms such as direct feedback alignment [Lillicrap et al., 2016, Nøkland, 2016], surrogate gradient learning Neftci et al. [2019] and many more. One must remain aware, however, that adding decorrelation to neural network architectures requires an additional weight matrix per neural network layer and induces additional computational overheads.
@@
Outside of theoretical treatments and application spheres, decorrelation has been presented in neuroscientific work as an explanation of the filtering which happens at multiple levels of nervous systems. The center-surround processing which takes places at the earliest stage of visual processing (at retinal ganglion cells), for example, has been proposed as an aid for decorrelation of visual input to the brain [Pitkow & Meister, 2012]. At a more general scale, inhibitory plasticity appears to be learned in a manner which also leads to spatial decorrelation [He & Cline, 2019], something which has also been modelled [Vogels et al., 2011]. Thus it appears that decorrelation might be active in real nervous systems and could therefore aid in whatever form of optimization is taking place.
@@
Beyond this, simulation work in computational neuroscience has also shown that via decorrelation of unit activities, competitive learning can be established to learn subspaces and carry out unsupervised feature extraction in a distributed and ‘local’ fashion [Bell & Sejnowski, 1997, Földiák, 1990, Oja, 1989, Zylberberg et al., 2011]. Thus, it appears that decorrelational processes may also have utility unsupervised competitive learning approaches.
However, in real nervous systems nearby neurons can have significantly correlated activities - a feature which may be required for redundancy and robustness to cell death. Thus, as strict a decorrelational process as presented herein seems unlikely. Nonetheless, we point toward a promising method by which local and distributed learning can be enabled, and it remains to be investigated as to how this could be combined mapped more directly to real neural systems.
We find that our decorrelation approach has a combination of benefits: increased convergence speed (per epoch) along with increased generalization performance, most notably for BP. One question which we address only shortly in this work is the relative tradeoff of decorrelation vs whitening processes. Wadia et al. [2021] demonstrate that whitening approaches must be regularized to maintain generalization performance. Given our results, we propose that decorrelation might be a sensible tradeoff, where signal correlations are removed but remaining signal (or noise) is not excessively rescaled.
@@ 6 Conclusion
In this work we present an integration of a set of research directions including natural gradient descent, decorrelation and whitening, as well as approximate methods for gradient descent. Notably we illustrate and describe how one component of natural gradients is often overlooked and can be framed as correlations in feature data (at all layers of a neural network) bringing parameters into a non-orthonormal relationship. We show that data correlation at every layer of a neural network can be removed, in a similar fashion to neural competition, to enable orders of magnitude faster training. These results and insights together suggest that failures of ‘biologically-plausible’ learning approaches and learning rules for distributed computing can be overcome through decorrelation and the return of parameters to an orthogonal basis.
@@ Appendix C A note on decorrelation vs whitening
In this work we limit our emphasis on whitening and focus more upon decorrelation. The reason for this is two-fold. First, the off-diagonal elements of the decorrelation matrices cause the greatest impact in skewing gradient descent. The diagonal elements simply scale up/down the gradient vector and this can be trivially dealt with by modern optimizers (e.g. with the Adam optimizer [Kingma & Ba, 2014]). Second, for data in which features are often zero or extremely sparse, normalizing for unit variance can result in extremely large valued activations (distributions with extremely long tails) or unstable updates in the stochastic updating regime.
@@ Appendix E Training pseudocode
Algorithm 1 describes pseudocode for the full computation and updating of a network which is carried out in this work. This is shown assuming the presentation of a single sample with its corresponding network update, though is in-practice used with a mini-batch of size 256 by default. The algorithm is shown for the case of backpropagation, though it is compatible with alternative credit assignment methods such as feedback alignment or node perturbation. Furthermore, the pseudocode shows the forward parameters as being updated by SGD, though in practice these are updated with the Adam [Kingma & Ba, 2014] optimizer.